OpenStudy (anonymous):

find the equation of the line thst passes through the points (4,5) and has a slope of -2

OpenStudy (jdoe0001):

\(\large {slope=m=-2\qquad \qquad \begin{array}{lllll} &x_1&y_1\\ &(4\quad ,&5) \end{array} \\\quad \\ y-y_1=m(x-x_1)}\)

OpenStudy (texaschic101):

just sub your stuff into the formula... y - 5 = -2(x - 4) can you finish from here ?

OpenStudy (texaschic101):

first distribute through the parenthesis and what do you get ?